A piece of ice of mass mice = 80g, which is initially at -15℃ is dropped into an isolated container with negligible heat capacity. The container contains 150g of water at 80℃. Determine the final temperature of the system? 


(Ci = 2093 J/kg℃, Cw = 4186 J/kg℃, Lf = 333 KJ/kg)


1
Expert's answer
2021-05-03T10:31:16-0400

Heat require to reach to temperature of ice at "0 ^oc"

"Q= m s\Delta T"

"=0.080\times 2093\times (15)"

"=2511.6J"

Heat require to melt the ice

"Q_1=mL"

"=0.080\times 333000 J"

"=26640J"

Hence, total energy require to melt the ice "=Q+Q_1"

"=26640+2511.6"

"=29151.6J"

Thermal energy of 150g of water at "80^\circ"

"=0.150\times 4200\times 80J"

"=50400J"

Hence we can say that ice will get melt completely.

"\Rightarrow 29151.6J=0.08\times 4200(t)+0.15\times 4200 \times (80-t)"

"\Rightarrow 336t+50400-630t=29151.6J"

"\Rightarrow 294t=21248.4J"


"\Rightarrow t=\frac{21248.4}{294}^\circ"

"\Rightarrow t= 72,27^\circ C"